# Quantitative BMO–BLO Estimates for the Hardy–Littlewood Maximal Function

### v1

@@ Abstract.
In this note, we study a quantitative extension of the John-Nirenberg inequality for the Hardy-Littlewood maximal function of a $\operatorname{BMO}$ function. More precisely, for every nonconstant locally integrable function $f$ such that MfMf is not identically infinite, we prove the inequality
@@
This research is supported by the Basque Government through the BERC 2022-2025 program, by the Ministry of Science and Innovation: Grant PRE2021-099091 funded by BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN / AEI / 10.13039/501100011033 and by ESF+.
@@ 1. Introduction and statement of the main results
for every function $f\in\operatorname{BMO}$ such that MfMf is not identically infinite. Introduced by Coifman and Rochberg in [CR], the space $\operatorname{BLO}$ consists of all locally integrable functions such that
@@
Theorem $1.$
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite, and let $w\in A_{\infty}$ be a weight. Then, for any cube $Q$ and $1\leq p<\infty$, we have
@@
The assumption that MfMf is not identically infinite cannot be removed. In [BDS], it is shown that this condition is necessary for the validity of the boundedness of $M$ on $\operatorname{BMO}$, and an explicit counterexample is provided by the function $f(x)=\log|x|$. We also refer to the recent work [BGSKM], which characterizes the class of measures $\mu$ for which $M\mu$ is finite almost everywhere.
@@
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. Let $w\in A_{\infty}$ be a weight.
@@
We can also prove that the converse implication holds: if the above inequality is satisfied for all locally integrable functions $f$ for which MfMf is not identically infinite, then the weight $w$ necessarily belongs to the class $A_{\infty}$. This equivalence yields a new quantitative characterization of the $A_{\infty}$ class in terms of the boundedness of the Hardy-Littlewood maximal operator acting on $\operatorname{BMO}$.
@@
where the supremum is taken over all functions $f$ such that $\|f\|_{\operatorname{BMO}}=1$ and MfMf is not identically infinite. As a consequence, we have the following quantitative equivalence of the Fujii-Wilson $A_{\infty}$ constant,
@@
  • Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. If $w\in A_{\infty}$, then for any cube $Q$ and $1\leq p<\infty$, we have
@@
    holds for every nonconstant locally integrable function $f$ such that MfMf is not identically infinite, every $1\leq p<\infty$, and every cube Q. Then $w\in A_{\infty}$ and $w]_{A_{\infty}}\lesssim_{n}C$.
@@ 2. The unweighted case
for each cube $Q$, and each $1\leq p<\infty$ (we refer to the proof of Theorem 1.1 below for more details). With this preliminary estimate established, we are now in a position to prove the unweighted version of Theorem $1.$
Proposition $2.$
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. Then, there exists a dimensional constant $c_{n}>0$ such that for every $1\leq p<\infty$, we have
@@
In the proof, we decompose MfMf into two terms: the local and nonlocal parts. The local contribution is estimated using (2.2), whereas the nonlocal term is treated following ideas from [Bennett].
@@
To estimate the second term, we bound MgMg pointwise in $Q$. Let $x\in Q$ and $P$ be a cube such that $x\in P$. We consider two cases. First, if $P\subseteq 3Q$, then
@@
Let $f$ be a nonconstant locally integrable function such that MfMf is not identically infinite. Then, there exists a dimensional constant $c_{n}>0$ such that
@@ 3. Weighted estimates and the characterization of $A_{\infty}$
where the supremum is taken over all functions $f$ such that $\|f\|_{\operatorname{BMO}}=1$ and MfMf is not identically infinite. We have to prove $w]_{A_{\infty}}\lesssim_{n}X$. Without loss of generality, we can assume that $X<\infty$.
@@
We can now estimate the $\operatorname{BLO}_{w}$ seminorm of MfMf:
@@
The first part follows directly from Theorem $1.$ Let us now prove the second part. Let $w$ be a weight and assume that inequality (1.8) holds for every nonconstant locally integrable function $f$ such that MfMf is not identically infinite and $\|f\|_{\operatorname{BMO}}=1$. In particular, if $f\in\operatorname{BMO}$, then taking $p=1$ in (1.8) we have

### v2

@@ Abstract.
In this note, we study a quantitative extension of the John-Nirenberg inequality for the Hardy-Littlewood maximal function of a $\operatorname{BMO}$ function. More precisely, for every nonconstant locally integrable function $f$ such that $Mf$ is not identically infinite, we prove the inequality
@@
This research is supported by the Basque Government through the BERC 2022-2025 program, by the Ministry of Science and Innovation through Grant PRE2021-099091 funded by BCAM Severo Ochoa accreditation CEX2021-001142-S/MICIN/AEI/10.13039/501100011033 and by ESF+, and by the project PID2023-146646NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ESF+.
@@ 1. Introduction and statement of the main results
for every function $f\in\operatorname{BMO}$ such that $Mf$ is not identically infinite. Introduced by Coifman and Rochberg in [CR], the space $\operatorname{BLO}$ consists of all locally integrable functions such that
@@
Theorem 1.1.
Let $f$ be a nonconstant locally integrable function such that $Mf$ is not identically infinite, and let $w\in A_{\infty}$ be a weight. Then, for any cube $Q$ and $1\leq p<\infty$, we have
@@
The assumption that $Mf$ is not identically infinite cannot be removed. In [BDS], it is shown that this condition is necessary for the validity of the boundedness of $M$ on $\operatorname{BMO}$, and an explicit counterexample is provided by the function $f(x)=\log|x|$. We also refer to the recent work [BGSKM], which characterizes the class of measures $\mu$ for which $M\mu$ is finite almost everywhere.
@@
Let $f$ be a nonconstant locally integrable function such that $Mf$ is not identically infinite. Let $w\in A_{\infty}$ be a weight.
@@
We can also prove that the converse implication holds: if the above inequality is satisfied for all locally integrable functions $f$ for which $Mf$ is not identically infinite, then the weight $w$ necessarily belongs to the class $A_{\infty}$. This equivalence yields a new quantitative characterization of the $A_{\infty}$ class in terms of the boundedness of the Hardy-Littlewood maximal operator acting on $\operatorname{BMO}$.
@@
where the supremum is taken over all functions $f$ such that $\|f\|_{\operatorname{BMO}}=1$ and $Mf$ is not identically infinite. As a consequence, we have the following quantitative equivalence of the Fujii-Wilson $A_{\infty}$ constant,
@@
  • Let $f$ be a nonconstant locally integrable function such that $Mf$ is not identically infinite. If $w\in A_{\infty}$, then for any cube $Q$ and $1\leq p<\infty$, we have
@@
    holds for every nonconstant locally integrable function $f$ such that $Mf$ is not identically infinite, every $1\leq p<\infty$, and every cube Q. Then $w\in A_{\infty}$ and $w]_{A_{\infty}}\lesssim_{n}C$.
@@ 2. The unweighted case
for each cube $Q$, and each $1\leq p<\infty$ (we refer to the proof of Theorem 1.1 below for more details). With this preliminary estimate established, we are now in a position to prove the unweighted version of Theorem 1.1.
Proposition 2.2.
Let $f$ be a nonconstant locally integrable function such that $Mf$ is not identically infinite. Then, there exists a dimensional constant $c_{n}>0$ such that for every $1\leq p<\infty$, we have
@@
In the proof, we decompose $Mf$ into two terms: the local and nonlocal parts. The local contribution is estimated using (2.2), whereas the nonlocal term is treated following ideas from [Bennett].
@@
To estimate the second term, we bound $Mg$ pointwise in $Q$. Let $x\in Q$ and $P$ be a cube such that $x\in P$. We consider two cases. First, if $P\subseteq 3Q$, then
@@
Let $f$ be a nonconstant locally integrable function such that $Mf$ is not identically infinite. Then, there exists a dimensional constant $c_{n}>0$ such that
@@ 3. Weighted estimates and the characterization of $A_{\infty}$
where the supremum is taken over all functions $f$ such that $\|f\|_{\operatorname{BMO}}=1$ and $Mf$ is not identically infinite. We have to prove $w]_{A_{\infty}}\lesssim_{n}X$. Without loss of generality, we can assume that $X<\infty$.
@@
We can now estimate the $\operatorname{BLO}_{w}$ seminorm of $Mf$:
@@
The first part follows directly from Theorem 1.1. Let us now prove the second part. Let $w$ be a weight and assume that inequality (1.8) holds for every nonconstant locally integrable function $f$ such that $Mf$ is not identically infinite and $\|f\|_{\operatorname{BMO}}=1$. In particular, if $f\in\operatorname{BMO}$, then taking $p=1$ in (1.8) we have